# Self-supervised Learning for Geospatial AI: A Survey

### v1

@@ IV Polylines
Polylines serve as an important spatial data type, widely employed across diverse geospatial data mining applications to depict a range of features, such as contour lines, road segments, and trajectories. SSL for polylines aims to derive effective representations for such data instances, with a particular emphasis on road segments and trajectories due to their prominence in existing literature. We will discuss these two representative instances of polylines in a geospatial context, and present the specialized SSL techniques tailored to each data instance. The surveyed studies for road networks are listed in Table II
IV-A Road Networks
IV-A1 Intrinsic Attributes
@@
Road networks are composed of connected road segments, where each segment is characterized as a polyline. As a collection of polylines, road networks are usually concepturalized as a graph. In this regard, two primary graph-based perspectives are adopted in the research. In the first perspective, road segments themselves are treated as nodes, and the connections between these segments are treated as graph edges. In the alternative perspective, road intersections serve as nodes, while the road segments linking these intersections are treated as edges.
@@
TABLE II: A summary of the surveyed papers on self-supervised learning for road networks.
@@
R2vec[77] predicative Random walk sampling None Skipgram Context neighbor prediction
IRN2Vec [79] predicative Random walk sampling Road attributes Cross-Entropy Road attribute prediction
RFN [81] predicative None Road attributes Cross-Entropy Graph reconstruction
RN2Vec [84] predicative Random walk sampling
@@
HyperRoad [82] predicative Hypergraph construction Road attributes
@@
HNRN [86] predicative None Trajectory Cross-Entropy Graph reconstruction
Toast [87] predicative
@@
Random walk sampling
@@
Random walk sampling
@@
TrajRNE [89] predicative Transition view construction Trajectory Cross-Entropy, MAE
@@
USPM [91]
predicative,
contrastive
None SVI
@@ IV-A2 Context Information
Apart from the topological structure and the attributes associated with road segments and intersections, road networks contain rich context information that can be leveraged to enhance the extraction of semantic knowledge. We introduce three types of context information utilized in existing methods, where road segments are mainly regarded as graph nodes.
@@
Spatial Features. Different from conventional graphs that primarily focus on connectivity, road networks are inherently defined within a geospatial context, exhibiting spatial features such as coordinates, lengths, and angles which are beneficial in modeling efforts [81, 82]. Accordingly, spatial information has been effectively utilized as context knowledge in various methods. RN2Vec [84] extends IRN2Vec to obtain representations for both intersections and road segments by selecting spatially nearby positive pairs among roads and intersections in skip-gram training. SARN [85] constructs a weighted adjacency matrix that reflects the spatial proximity among road segments. The matrix is inferred from both road connectivity and the spatial and angular distances between road segments. Utilzing this matrix, SARN employs a GNN-based contrastive learning approach [93] with graph augmentation techniques, deriving similar representations for the identical road segments from two distinct augmented graphs.
@@
Trajectories. Trajectories traveled on road networks act as valuable data sources that provide travel-related semantics beyond the topological structure. HNRN [86] develops a hierarchical GNN framework to model relationships from individual road segments to structural regions and further extending to functional zones. It involves incorporating trajectories to derive connectivity matrix in structural regions, and aims to reconstruct the connectivity matrix for the base-level road segments derived from the higher level of region and functional representations. Toast [87] is the initial effort to explicitly utilize detailed trajectories to enhance road network representation learning through predicative SSL. It equips skip-gram training with additional traffic and attribute prediction for context neighbors. Besides, it proposes two novel trajectory pre-training tasks within BERT framework [38]: route recovery, which recovers a sequence of masked road segments, and trajectory discrimination, which assesses whether a trajectory is authentic or simulated. These tasks enable the encoding of transition patterns and long-term dependencies intrinsic in road networks. DyToast [88] further extends this method by incorporating temporal consideration into the representations. It modifies the skip-gram training and the BERT framework by integrating parameterized trigonometric functions to capture dynamics and evolution in time dimension. JCLRNT [90] proposes three types of objectives based on contrastive SSL: road-road, trajectory-trajectory and road-trajectory contrastive loss. Each type is designed to differentiate entity pairs that share relatedness (e.g., road segments frequently traveled within trajectories) with those that do not. TrajRNE [89] utilizes the road transition matrix derived from trajectories in GNN aggregation function, and employs the objective of reconstructing the original topological structure of road networks from this transition matrix.
@@ IV-A3 Applications
The representations of road networks, derived from either predicative or contrastive SSL objectives, provide task-agnostic inputs that can be directly utilized or fine-tuned with geospatial encoder, for downstream applications. For example, they are directly applicable in classifying attributes of road segments or intersections with simple models (e.g., MLP), such as road type [89], lane number [82], and intersection tags [79, 84]. Besides, these representations are similarly utilized to infer the traffic status of road segments, including metrics like average speed [87, 90] and speed limits [81]. They also support the efficient vector computation for road network-based queries, such as calculating shortest path distances [85]. Furthermore, road segment representations serve as the effective start point for training in models that involve map-matched trajectories, for applications such as destination prediction [86] and anomalous sub-trajectory detection [94]. By enabling a diverse range of analytical tasks related to road networks, these representations offer substantial potential to improve understanding and decision-making within the domain of transportation infrastructure.
IV-B Trajectory
IV-B1 Intrinsic Attributes
A trajectory, composed of a sequence of sampled points that represent the path of a moving object, can essentially be conceptualized as a polyline. Based on the definition, the context feature associated with each point in trajectory data instance includes timestamps, along with potentially additional textual content, and semantic tags, etc. The objective of SSL for trajectory data is to learn a representation produced by geospatial encoder $f_{\theta}$ for any given trajetocry $\mathcal{T}$: $\mathbf{e}_{T}=f_{\theta}(\mathcal{T})\in\mathbb{R}^{d}$. The sequential nature of a trajectory, marked by its spatio-temporal point sequence, represents its most fundamental intrinsic attributes. Accordingly, sequential models are used to model trajectory data, effectively capturing the transition patterns and long-term dependencies inherent in this data instance. The surveyed studies for road networks are listed in Table III
In the category of predicative SSL methods, trajectories are typically trained to reconstruct their original sequence from a corrupted version of the input. For example, TCDRL [95] proposes to extract local features within sliding windows applied on trajectory records, such as speeds and rate of turns. These features are then processed using a sequence-to-sequence encoder-decoder architecture [96] based on RNN to reconstruct the local features for each window. t2vec [97] refines this reconstruction process by aligning with the tokenization paradigm in natural language. It partitions the geospatial space into regular and square-sized grids and match the coordinates to their corresponding grids (tokens). Moreover, t2vec introduces downsampling and point distortion in the input sequence, and aims to reconstruct the original trajectory with spatial proximity aware loss that penalizes more for the predicted tokens that deviate significantly from the correct grids. In such an encoder-decoder framework, the vector produced by encoder part can be treated as the trajectory representation.
This framework is enhanced by integrating a variety of modifications in terms of model architecture and loss functions. Specifically, model adaptations include refinements via variatioanl inference [98] or self-attention mechanisms [99]. Geo-Tokenizer [100] reduces the number of grids to be trained by representing a location as a combination of multiple shared grids at several granular scales. It then utilizes the objective of predicting the grids for the next token at each scale in a hierarchical way. Furthermore, AdvTraj2vec [101] aims to learn more robust trajectory representations through adversarial training. It adds perturbations to the token embeddings of the input sequence, with the magnitude of these perturbations guided by generative adversarial network (GAN) [102] to ensure the effects are neither too large or too small. E2DTC incorporates self-training via soft cluster assignments [103] as an auxiliary loss into the reconstruction process. STPT [104] performs a sub-trajectory discrimination loss that differentiates whether pairs of sub-trajectory representations originate from the same source.
In addition to considering spatial dimension associated with trajectories, several SSL methods leverage temporal dimension into the reconstruction process. The method in [105] expands the square-sized spatial partitions to 3D spatio-temporal grids for each point. It accordingly adapts the reconstruction loss to account for temporal effects, imposing heavier penalties for reconstructed results with larger temporal discrepancies. Similarly, RSTS [106] also employs 3D spatio-temporal grids and applies a linear combination of spatial and temporal distances for reconstructed results as the final loss. Besides, DeepTEA [107] utilizes Convolutional-LSTM [108] to model historical traffic conditions reflected in holistic trajectories, thus enhancing the variational inference by providing additional hints into dynamic patterns for each trajectory.
@@ IV-B1 Intrinsic Attributes
TABLE III: A summary of the surveyed papers on self-supervised learning for Trajectory.
@@
TCDRL[95] predicative
Sliding window feature extraction
@@
GM-VSAE [98] predicative None Grid partition Cross-Entropy Variational trajectory reconstruction
Geo-Tokenizer [100] predicative None Multi-scale grid partition Multi-scale Cross-Entropy
Autoregressive next grid prediction
AdvTraj2vec [101] predicative
@@
Spatial-aware Cross-Entropy,
@@
Spatial-aware Cross-Entropy,
@@
STPT [104] predicative None
@@
3D temporal grid partition
@@
3D temporal grid partition
Spatial-aware Cross-Entropy,
@@
DeepTEA [107] predicative None
@@
Variational trajectory reconstruction
@@
Trajectory trimming&simplification
@@
CSSRNN [84] predicative Map matching Road networks Cross-Entropy Autoregressive road prediction
Traj2Vec [112] predicative Map matching Road networks Cross-Entropy, MSE
@@
Road-road, Road-trajectory contrast
@@
CTLE [121] predicative
Location masking,
@@
At2vec [122] predicative
Location dropping
@@
Spatial-temporal-activity-aware
@@
Location embedding perturbation,
Latent space perturbation
@@
IV-B2 Context Information
In addition to the spatial and temporal attributes intrinsic to trajectories, we discuss two distinct scenarios where trajectories are modeled under specific constraints and characteristics with further context information.
@@ IV-B2 Context Information
Several predicative SSL methods for map-matched trajectories utilize objectives similar to those used in trajectories with grid partitions, such as the reconstruction of road segments [112, 113] and next road prediction [127]. Besides, tailored objectives are devised to effectively incorporate the knowledge within road networks. For example, JGRM [114] aims to recover the road segments masked from the complete path. It also considers the original point sequence and adopts another objective of differentiating whether pairs of representations are derived from the same trajectory with both the road segment sequence and the point sequence.
@@
For contrastive SSL methods applied to map-matched trajectories, ST2Vec [128] adopts co-attention mechanism to merge spatial and temporal embeddings, followed by LSTM sequence modeling. This model employs energy-based margin functions (i.e., triplet loss) to enforce higher similarities for positive pairs which follow similar routes. GRLSTM [115] combines GNN and graph embedding techniques [129] to handle sequence inputs and augments LSTM with residual connections to generate trajectory representations. It employs similar triplet loss at both the point level and the trajectory level. In addition, PIM [116] treats trajectories that share the same source and destination as positive pairs, and aims to maximize mutual information over these pairs. MMTEC [117] utilizes both a discrete sequence model based on Transformer, and a continuous model formulated as neural controlled differential equation to generate representations from two views. It applies a different contrastive learning objective of maximizing entropy coding between two views.
@@
Moreover, both contrastive and predictive SSL can be simultaneously employed with a single framework. HMTRL [118] employs the strategies to predict road attributes and traverse time for road segments, as well as employing full trajectory and its sub-trajectory for contrastive learning. START [119] incorporates trajectory pattern into GNN aggregation and enhances Transformer with time-aware self-attention mechanism. Furthermore, it utilizes masked road segment recovery as the predicative task, and enhances its contrastive learning aspect by employing four distinct data augmentation operation –trajectory trimming, masking, temporal shifting, and dropout– to generate positive pairs. Similarly, LightPath [120] utilizes masked road path recovery as its predictive task. For its contrastive aspect, LightPath achieves pairwise matching through the use of dual encoders and varied dropping ratios to generate positive pairs from the same trajectories.
@@
Semantic Information. Trajectories may carry rich semantic meaning when composed of check-in sequence at specific POIs, which provide concrete location names and related activities (i.e., categories) but result in sparser sequences. Several SSL methods are developed to tackle semantic trajectories. CTLE [121] designs two predicative tasks, namely masked location recovery and masked hour prediction, with a sinusoidal temporal encoding technique incorporated into Transformer to derive representations. At2vec [130, 122] utilizes an encoder-decoder framework for sequence reconstruction, and adopts multi-level attention to consider the importance of semantic information at different locations. SML-LP [123] applies trajectory augmentation techniques that modify several locations within close spatial and temporal proximity to form positive pairs, and aims to maximize the mutual information between the LSTM-derived hidden representations and the future location representations. Furthermore, CACSR [124] innovates by generating challenging positive and negative pairs in the representation space, rather than input sequence, via adversarial perturbations. We note that while SSL techniques are integrated into the modeling of semantic trajectories in other studies [131, 132], the main target and its objective derives from supervised applications. Therefore, these specific studies are not further discussed in our context.
IV-B3 Applications

### v2

@@ 4 Polylines
Polylines serve as an important spatial data type, widely employed across diverse GeoAI applications to depict a range of features, such as contour lines, road segments, and trajectories. SSL for polylines aims to derive effective representations for such data instances, with a particular emphasis on road segments and trajectories due to their prominence in existing literature. While both road networks and trajectories can be geometrically represented as polylines, they differ fundamentally in structure and semantics. Road networks typically denote static infrastructure with topological consistency and connectivity, whereas trajectories capture dynamic movements, often annotated with timestamps and user-specific metadata. This divergence in semantics and application motivates a distinct categorization for these two data instances in our survey, enabling a more precise analysis of SSL methods tailored to each context. Consequently, we will separately discuss these two instances of polylines, and present the specialized SSL techniques tailored to each data instance. The surveyed studies for road networks are listed in Table 2.
4.1 Road Networks
4.1.1 Intrinsic Attributes
@@
Road networks are composed of connected road segments, each of which is represented as a polyline. Since the fundamental analytical units in road network studies are typically individual road segments rather than enclosed regions, road networks are classified under the category of polylines. Conceptually, road networks can be represented as graphs, allowing for the modeling and analysis of their structural and topological properties. In this regard, two primary graph-based perspectives are adopted in the research. In the first perspective, road segments themselves are treated as nodes, and the connections between these segments are treated as graph edges. In the alternative perspective, road intersections serve as nodes, while the road segments linking these intersections are treated as edges.
@@
Table 2: A summary of the surveyed papers on self-supervised learning for road networks.
@@
R2vec[77] Predictive Random walk sampling None Skipgram Context neighbor prediction
IRN2Vec [78] Predictive Random walk sampling Road attributes Cross-Entropy Road attribute prediction
RFN [79] Predictive None Road attributes Cross-Entropy Graph reconstruction
RN2Vec [81] Predictive Random walk sampling
@@
HyperRoad [80] Predictive Hypergraph construction Road attributes
@@
HNRN [83] Predictive None Trajectory Cross-Entropy Graph reconstruction
Toast [84] Predictive
@@
Random walk sampling,
Similar trajectory generation
@@
Random walk sampling,
Similar trajectory generation
@@
TRACK [88] Predictive, Contrastive
Road segment masking,
Traffic state masking,
Similar trajectory generation
Trajectory
Cross-Entropy, MAE,
InfoNCE
Time-aware masked road&
traffic state recovery,
Traffic state-road matching,
Trajectory discrimination
TrajRNE [86] Predictive Transition view construction Trajectory Cross-Entropy, MAE
@@
USPM [90] Predictive, Contrastive None SVI
@@ 4.1.2 Context Information
Apart from the topological structure and the attributes associated with road segments and intersections, road networks contain rich context information that can be leveraged to enhance the extraction of semantic knowledge. We introduce several types of context information utilized in existing methods, where road segments are mainly regarded as graph nodes.
@@
Spatial Attributes. Different from conventional graphs that primarily focus on connectivity, road networks are inherently defined within a geospatial context, exhibiting spatial features such as coordinates, lengths, and angles which are beneficial in modeling efforts [79, 80]. Accordingly, spatial information has been effectively utilized as context knowledge in various methods. RN2Vec [81] extends IRN2Vec to obtain representations for both intersections and road segments by selecting spatially nearby positive pairs among roads and intersections in skip-gram training. SARN [82] constructs a weighted adjacency matrix that reflects the spatial proximity among road segments. The matrix is inferred from both road connectivity and the spatial and angular distances between road segments. Utilzing this matrix, SARN employs a GNN-based contrastive learning approach [166] with graph augmentation techniques, deriving similar representations for the identical road segments from two distinct augmented graphs.
@@
Trajectories. Trajectories traveled on road networks act as valuable data sources that provide travel-related semantics beyond the topological structure. HNRN [83] develops a hierarchical GNN framework to model relationships from individual road segments to structural regions and further extending to functional zones. It involves incorporating trajectories to derive connectivity matrix in structural regions, and aims to reconstruct the connectivity matrix for the base-level road segments derived from the higher level of region and functional representations. Toast [84] is the initial effort to explicitly utilize detailed trajectories to enhance road network representation learning through predictive SSL. It equips skip-gram training with additional traffic and attribute prediction for context neighbors. Besides, it proposes two novel trajectory pre-training tasks within BERT framework [37]: route recovery, which recovers a sequence of masked road segments, and trajectory discrimination, which assesses whether a trajectory is authentic or simulated through random walk sampling on the road network graph. These tasks enable the encoding of transition patterns and long-term dependencies intrinsic in road networks. DyToast [85] further extends this method by incorporating temporal consideration into the representations. It modifies the skip-gram training and the BERT framework by integrating parameterized trigonometric functions to capture dynamics and evolution in time dimension. Similarly, TRACK [88] considers temporal dynamics for road segment representations by integrating trajectory data-based transition probabilities into the GAT attention mechanism while modeling traffic states for road networks. Building on Toast pretext tasks, TRACK further employs masked traffic state imputation and prediction tasks for training dynamic road segment representations. Besides, it further incorporates a co-attentional transformer encoder with a gravity-based attention mechanism and a trajectory-traffic state matching task, ensuring mutual reinforcement between these two data sources. JCLRNT [87] proposes three types of objectives based on contrastive SSL: road-road, trajectory-trajectory and road-trajectory contrastive loss. Each type is designed to differentiate entity pairs that share relatedness (e.g., road segments frequently traveled within trajectories) with those that do not. TrajRNE [86] utilizes the road transition matrix derived from trajectories in GNN aggregation function, and employs the objective of reconstructing the original topological structure of road networks from this transition matrix.
@@ 4.1.3 Applications
The representations of road networks, derived from either predictive or contrastive SSL objectives, provide task-agnostic inputs that can be directly utilized or fine-tuned with geospatial encoder, for downstream applications. For example, they are directly applicable in classifying attributes of road segments or intersections with simple models (e.g., MLP), such as road type [86], lane number [80], and intersection tags [78, 81]. Besides, these representations are similarly utilized to infer the traffic status of road segments, including metrics like average speed [84, 87] and speed limits [79]. They also support the efficient vector computation for road network-based queries, such as calculating shortest path distances [82]. Furthermore, road segment representations serve as the effective start point for training models that involve map-matched trajectories, for applications such as destination prediction [83] and anomalous sub-trajectory detection [167]. By enabling diverse analytical tasks related to road networks, these representations offer substantial potential to improve understanding and decision-making within the domain of transportation infrastructure.
4.2 Trajectory
4.2.1 Intrinsic Attributes
A trajectory, composed of a sequence of sampled points that represent the path of a moving object, can essentially be conceptualized as a polyline. Based on the definition, the context feature associated with each point in trajectory data instance includes timestamps, along with potentially additional textual
@@ 4.2.1 Intrinsic Attributes
Table 3: A summary of the surveyed papers on self-supervised learning for trajectory.
@@
TCDRL[91] Predictive
Sliding window
feature extraction
@@
GM-VSAE [93] Predictive None Grid partition Cross-Entropy Variational trajectory reconstruction
Geo-Tokenizer [94] Predictive None Multi-scale grid partition Multi-scale Cross-Entropy
Autoregressive next grid
prediction
AdvTraj2vec [95] Predictive
@@
Spatial-aware
Cross-Entropy,
@@
Spatial-aware
Cross-Entropy,
@@
STPT [97] Predictive None
@@
3D temporal
grid partition
@@
3D temporal
grid partition
Spatial-aware
Cross-Entropy,
@@
DeepTEA [99] Predictive None
@@
Variational trajectory
reconstruction
UniTraj [100] Predictive
Point dropping
Point coordinates
Cross-Entropy
Trajectory reconstruction
@@
Trajectory trimming
& simplification
@@
CSSRNN [105] Predictive Map matching Road networks Cross-Entropy Autoregressive road prediction
Traj2Vec [106] Predictive Map matching Road networks Cross-Entropy, MSE
@@
Road-road, Road-trajectory
contrast
@@
GREEN [116]
Predictive,
Contrastive
Map matching,
Road segment masking
Grid partition,
Road networks
Cross-Entropy,
InfoNCE
Masked road recovery,
Road-grid view trajectory contrast
CTLE [117] Predictive
Point masking,
@@
At2vec [118] Predictive
Point dropping
@@
Spatial-temporal
-activity-aware
@@
CLMTR [120] Contrastive
Point dropping,
Spatial distortion,
Trajectory trimming
& simplification
Location semantics
InfoNCE
Spatial-temporal
-textual feature contrast,
Trajectory-trajectory contrast
@@
Location embedding & latent
space perturbation
@@
MVTraj [124]
Predictive,
Contrastive
Map matching,
Road segment masking,
Point dropping
Grid partition,
Road networks,
Location semantics
Cross-Entropy,
InfoNCE
Masked grid/road recovery,
Multi-view trajectory
cross contrast
PTrajM [123]
Contrastive
Map matching,
Point allocation
Road networks,
Location semantics
InfoNCE
Trajectory-poi view contrast,
Trajectory-road view contrast
RED [122]
Predictive
Map matching
Road networks,
User information
Cross-Entropy
Masked road recovery,
Next segment prediction
MM-Path [125]
Predictive,
Contrastive
Map matching,
Road segment masking
Road networks,
Road images
Triple loss,
InfoNCE,
Cross-Entropy
Masked road recovery,
Multi-granularity road-image
view contrast,
Fused road-image view contrast
content, and semantic tags, etc. The objective of SSL for trajectory data is to learn a representation produced by geospatial encoder $f_{\theta}$ for any given trajetocry $\mathcal{T}$: $\mathbf{e}_{T}=f_{\theta}(\mathcal{T})\in\mathbb{R}^{d}$. The sequential nature of a trajectory, marked by its spatio-temporal point sequence, represents its most fundamental intrinsic attributes. Accordingly, sequential models are used to model trajectory data, effectively capturing the transition patterns and long-term dependencies inherent in this data instance. The surveyed studies for trajectory are listed in Table LABEL:tab:_trajectory.
For predictive SSL methods, trajectories are typically trained to reconstruct their original sequence from a corrupted version of the input. For example, TCDRL [91] proposes to extract local features within sliding windows applied on trajectory records, such as speeds and rate of turns. These features are then processed using a sequence-to-sequence encoder-decoder architecture [168] based on RNN to reconstruct the local features for each window. t2vec [92] refines this reconstruction process by aligning with the tokenization paradigm in natural language. It partitions the geospatial space into regular and square-sized grids and match the coordinates to their corresponding grids (tokens). Moreover, t2vec introduces downsampling and point distortion in the input sequence, and aims to reconstruct the original trajectory with spatial proximity aware loss that penalizes more for the predicted tokens that deviate significantly from the correct grids. In such an encoder-decoder framework, the vector produced by encoder part can be treated as the trajectory representation.
This framework is enhanced by integrating a variety of modifications in terms of model architecture and loss functions, such as integrating via variatioanl inference [93] or self-attention  [169]. Moreover, Geo-Tokenizer [94] reduces the number of grids to be trained by representing a location as a combination of multiple shared grids at several granular scales. It utilizes the objective of predicting the grids for the next token at each scale in a hierarchical way. Furthermore, AdvTraj2vec [95] aims to learn more robust trajectory representations through adversarial training. It adds perturbations to the token embeddings of the input sequence, with the magnitude of these perturbations guided by generative adversarial network [170] to ensure the effects are neither too large or too small. E2DTC incorporates self-training via soft cluster assignments [96] as an auxiliary loss into the reconstruction process. STPT [97] performs a sub-trajectory discrimination loss that differentiates whether pairs of sub-trajectory representations originate from the same source.
In addition to considering spatial dimension associated with trajectories, several SSL methods leverage temporal dimension into the reconstruction process. The method in [98] expands the square-sized spatial partitions to 3D spatio-temporal grids for each point. It accordingly adapts the reconstruction loss to account for temporal effects, imposing heavier penalties for reconstructed results with larger temporal discrepancies. Similarly, RSTS [104] also employs 3D spatio-temporal grids and applies a linear combination of spatial and temporal distances for reconstructed results as the final loss. Besides, DeepTEA [99] utilizes Convolutional-LSTM [171] to model historical traffic conditions reflected in holistic trajectories, thus enhancing the variational inference by providing additional hints into dynamic patterns for each trajectory. Recently, UniTraj [100] introduces a large-scale, global trajectory dataset to train a universal trajectory model capable of generalizing across diverse tasks and regions, thus utilizing pure spatial coordinates without partitioning. To manage variations in sampling rates, it employs dynamic resampling and utilizes multiple masking strategies, including random, block, key point, and last-N masking, for data augmentation. The model adopts an encoder-decoder architecture with Rotary Position Embedding [172] to model trajectory sequences, and aims to predict masked segments for training.
4.2.2 Context Information
In addition to the intrinsic spatial and temporal attributes, we discuss two distinct scenarios where trajectories are modeled under specific constraints and characteristics with further context information.
@@ 4.2.2 Context Information
Several predictive SSL methods for map-matched trajectories utilize objectives similar to those used in trajectories with grid partitions, such as the reconstruction of road segments [106, 107] and next road prediction [105]. Besides, tailored objectives are devised to effectively incorporate the knowledge within road networks. For example, JGRM [108] aims to recover the road segments masked from the complete path. It also considers the original point sequence and adopts another objective of differentiating whether pairs of representations are derived from the same trajectory with both the road segment sequence and the point sequence.
@@
For contrastive SSL methods applied to map-matched trajectories, ST2Vec [109] adopts co-attention mechanism to merge spatial and temporal embeddings, followed by LSTM sequence modeling. This model employs energy-based margin functions (i.e., triplet loss) to enforce higher similarities for positive pairs which follow similar routes. GRLSTM [110] combines GNN and graph embedding techniques [175] to handle sequence inputs and augments LSTM with residual connections to generate trajectory representations. It employs similar triplet loss at both the point level and the trajectory level. In addition, PIM [111] treats trajectories that share the same source and destination as positive pairs, and aims to maximize mutual information over these pairs. MMTEC [112] utilizes both a discrete sequence model based on Transformer, and a continuous model formulated as neural controlled differential equation to generate representations from two views. It applies a contrastive learning objective of maximizing entropy coding between two views.
@@
Moreover, both contrastive and predictive SSL can be simultaneously employed with a single framework. HMTRL [113] employs the strategies to predict road attributes and traverse time for road segments, as well as employing full trajectory and its sub-trajectory for contrastive learning. START [114] incorporates the trajectory pattern in GNN aggregation and enhances the Transformer with a time-sensitive self-attention mechanism. Furthermore, it utilizes masked road segment recovery as the predictive task, and enhances its contrastive learning aspect by employing four distinct data augmentation operation – trajectory trimming, masking, temporal shifting, and dropout– to generate positive pairs. Similarly, LightPath [115] utilizes masked road path recovery as its predictive task. For its contrastive aspect, LightPath achieves pairwise matching through the use of dual encoders and varied dropping ratios to generate positive pairs from the same trajectories. GREEN [116] integrates a grid encoder based on CNN and a road encoder based on GNN to encode trajectories from two perspectives. It employs contrastive training to align the representations from both encoders for the same trajectory while using masked road segment recovery as a predictive task.
@@
Semantic Information. Trajectories may carry rich semantic meaning when composed of check-in sequence at specific POIs, which provide concrete location names and related activities (i.e., categories) but result in sparser sequences. Several SSL methods are developed to tackle semantic trajectories. CTLE [117] designs two predictive tasks, namely masked location recovery and masked hour prediction, with a sinusoidal temporal encoding technique incorporated into Transformer to derive representations. At2vec [176, 118] utilizes an encoder-decoder framework for sequence reconstruction, and adopts multi-level attention to consider the importance of semantic information at different locations. SML-LP [119] applies trajectory augmentation techniques that modify several locations within close spatial and temporal proximity to form positive pairs, and aims to maximize the mutual information between the LSTM-derived hidden representations and the future location representations. CLMTR [120] encodes trajectory by leveraging textual descriptions of POIs processed with BERT alongside spatial and temporal embeddings to capture geographical proximity and periodicity. It employs contrastive learning at two levels: intra-trajectory, where textual features are contrasted with fused spatiotemporal features within the same trajectory, and inter-trajectory, where nearest-neighbor trajectories serve as positive pairs while distant ones act as negatives. Furthermore, CACSR [121] innovates by generating challenging positive and negative pairs in the representation space, rather than input sequence, via adversarial perturbations. We note that while SSL techniques are integrated into the modeling of semantic trajectories in other studies [177, 178], the main target and its objective derives from supervised applications. Therefore, these specific studies are not further discussed in our context.
Multiple Context Sources. The extensive studies have revealed that different types of context information provide distinct and complementary insights for trajectory data representation. For instance, spatial coordinates and grid partitions offer high-resolution positional accuracy, while road networks impose movement constraints and reveal structural mobility patterns. Additionally, location semantics provide valuable information regarding regional functionalities and human activity patterns. To this end, recent research has increasingly focused on integrating multiple context sources simultaneously, recognizing that their combination enables a more comprehensive and multi-dimensional understanding of trajectory data. MVTraj [124] integrates context information from multiple views, including GPS trajectories, road networks and POIs. It employs three encoders to capture point sequences, road network paths, and grid-based semantic features, utilizing a hierarchical multi-modal interaction with twelve attention streams for cross-view fusion. The model incorporates masked part recovery as a predictive task and applies contrastive learning to align trajectory representations across different views. PTrajM [123] utilizes the Mamba state-space model [179] to capture continuous movement behaviors and infer travel purposes from road network and POI semantic views by assigning data points to POIs. It extracts road-based and POI-based trajectory representations using pre-trained textual embeddings and Transformer-based encoders. A contrastive learning approach is then employed to align trajectory representations extracted from the state-space model with their corresponding road network and POI-based representations. Several studies have expanded trajectory representation learning by integrating additional contextual information. RED [122] incorporates user information alongside spatial and temporal encodings to enhance trajectory representation within a road network context. It employs dual Transformer models for next-segment prediction and masked road segment prediction, adapting the latter to selectively mask non-crucial path segments as a pretext task. MM-Path [125] combines road network-based trajectories with corresponding road segment images across three granularities: road intersection level, road sub-path level, and entire path level. It employs Transformer-based encoders for both modalities and introduces a GNN-based cross-modal fusion to integrate image and road network representations. The model utilizes masked road segment recovery as a predictive task and applies contrastive learning by aligning representations across modalities and granularities.
4.2.3 Applications